OpenStudy (anonymous):

write in radical notation 81^-(3/4) then evaluate the radical expression

OpenStudy (anonymous):

ma'am is the question complete ?

OpenStudy (ajprincess):

\(\large(a^b)^c=a^{b*c}\) \(\large a^{-b}=\frac{1}{a^b}\) \(81=3*3*3*3*3=3^4\) Using these can u do ur question @mathgirl501?

OpenStudy (abb0t):

Aj is correct, but just to clarify things a bit more: \[(a^b)^{\frac{ 1 }{ c }} = a^{\frac{ b }{ c }} = \sqrt[c]{a^b}\]